# ExMAG: Learning of Maximally Ancestral Graphs

### v1

@@ 4.2 MILP Formulation
The cost function for the Mixed Integer Quadratic Program of ExMAG is the following $l_{p}$ norm:
@@
Lastly, we must enforce conditions for directed or almost directed cycles and inducing paths. Those conditions are enforced lazily using a separation routine explained later. Directed cycles are enforced in a way adopted from [38]. Therefore, they are left out of this paper. Suppose an almost directed cycle is formed by edges in set $E$ and a bidirected edge $(x,y)$. Then, this almost directed cycle is forbidden by constraint
$b_{x,y}+\sum_{(j,k)\in E}e_{j,k}\leq|E|.$
@@ 6.1 Used Datasets
The second dataset uses randomly generated bow-free (BF) graphs, a subset of all possible MAGs. A bow-free graph is defined as a graph that does not contain a pair of vertices $i,j$, such that $i\rightarrow j$ and at the same time $i\leftrightarrow j$. The BF graph generation process has two parameters. First, the probability of a directed edge. Second, the probability of bidirected edge. The generation process is as follows. First, a bow-free graph with the given edge probabilities is generated randomly. The weights of the sampled graphs are randomly sampled from the set $(-2.0,-0.5)\cup(0.5,2.0)$. The adjecancy matrix of the directed edges defines the weights of the structural equation model. Then the data samples are generated using the structural equation, where the noise is sampled from multivariete Gaussian distribution with covariance matrix equals to the adjecancy matrix of bidirected edge generated in the previous step.
@@ 6.2 Evaluation Criteria
Suppose that a tested algorithm produced weight matrix $W_{\cdot}$. Such a matrix can contain nearly zero weights. For such reasons, thresholding is done, keeping only edges with weight greater or equal to $\delta$. In cases when the ground-truth weight matrix $W$ is known, the best solution (in terms of structural Hamming distance, see below) is kept over those defined by different threshold $\delta$ values.
@@ 6.4 Experimental Results
The results on the real-world dataset can be seen in 2. Contrary to the original expectations, the highest risk importer is company 2H66B7, which stands for Lufthansa. The second highest risk importer is 2H6677, i.e., the Deutsche Bank, which is an expected result.
@@ 7 Conclusion
This work has been partially supported by the CoDiet project, co-funded by the European Union under Horizon Europe grant number 101084642.
This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.

### v2

@@ 4.2 MILP Formulation
The cost function for the Mixed Integer Quadratic Program of ExMAG is the following $l_{q}$ norm:
@@
Lastly, we must enforce conditions for directed or almost directed cycles and inducing paths. Those conditions are enforced lazily using a separation routine explained later. Directed cycles are enforced in a way adopted from [46]. Therefore, they are left out of this paper. Suppose an almost directed cycle is formed by edges in set $E$ and a bidirected edge $(u,v)$. Then, this almost directed cycle is forbidden by constraint
$b_{u,v}+\sum_{(j,k)\in E}e_{j,k}\leq|E|.$
@@ 6.1 Used Datasets
The second dataset uses randomly generated bow-free (BF) graphs, a subset of all possible MAGs. A bow-free graph is defined as a graph that does not contain a pair of vertices $i,j$, such that $i\rightarrow j$ and at the same time $i\leftrightarrow j$. The BF graph generation process has two parameters. First, the probability of a directed edge. Second, the probability of bidirected edge. The generation process is as follows. First, a bow-free graph with the given edge probabilities is generated randomly. The weights of the sampled graphs are randomly sampled from the set $(-2.0,-0.5)\cup(0.5,2.0)$. The adjacency matrix of the directed edges defines the weights of the structural equation model. Then the data samples are generated using the structural equation, where the noise is sampled from multivariete Gaussian distribution with covariance matrix equals to the adjacency matrix of bidirected edge generated in the previous step.
@@ 6.2 Evaluation Criteria
Suppose that a tested algorithm produced weight matrix $\hat{W}$. Such a matrix can contain nearly zero weights. For such reasons, thresholding is done, keeping only edges with weight greater or equal to $\delta$. In cases when the ground-truth weight matrix $W$ is known, the best solution (in terms of structural Hamming distance, see below) is kept over those defined by different threshold $\delta$ values.
@@ 6.4 Experimental Results
The results on the real-world dataset can be seen in 3. Contrary to the original expectations, the highest risk importer is company 2H66B7, which stands for Lufthansa. The second highest risk importer is 2H6677, i.e., the Deutsche Bank, which is an expected result.
@@ 7 Conclusion
This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101084642. The work of G. K. has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
@@ Appendix B SHD Results
Appendix C Table of Notation
Table 1: A Table of Notation.
Symbol Representation
$G$ Gender
$D$ Department
$A$ Admit acceptance
$U$ potential confounder, is academic ability in this example
$\mathcal{G}$ Graph
$\mathcal{V}$ Vertex set of the graph
$\mathcal{E}$ Set of directed edges
$\mathcal{U}$ Set of undirected edges
$\text{pa}_{\mathcal{G}}(v)$ Parents of vertex $v$ in graph $\mathcal{G}$
$\text{sp}_{\mathcal{G}}(v)$ Spouses of vertex $v$ in graph $\mathcal{G}$
$\text{an}_{\mathcal{G}}(v)$ Ancestors of vertex $v$ in graph $\mathcal{G}$
$\text{dis}_{\mathcal{G}}(v)$ District of vertex $v$ in graph $\mathcal{G}$
$\text{collider}_{\mathcal{G}}(u)$ Collider nodes in graph $\mathcal{G}$
$W$ Weight matrix
$\hat{W}$ Tested weight matrix
$E$ Directed edge matrix
$B$ Bidirected edge matrix
$F$ Indicating Matrix (existence of a direct causal relationship)
$X_{i,j}$ Value of the $j$-th variable for the $i$-th data point
$w_{k,j}$ Weight of the edge from variable $k$ to variable $j$
$e_{j,k}$ Binary variable indicating a directed edge from $j$ to $k$
$b_{j,k}$ Binary variable indicating a bidirected edge between $j$ and $k$
$f_{j,k}$ Binary variable indicating a existence of directed edge between $j$ and $k$
$d_{j,k}$ Binary variable indicating a directed edge between $j$ and $k$
$r_{ij}$ Contribution of pair $(i,j)$ to SHD
$\lambda$ Regularization parameter
$c$ Large constant for weight bounding
$d$ distance matrix
$n$ Number of data points
$q$ Exponent in the cost function ($q=1$ or $q=2$)
$GT$ Edge type in the ground truth graph
$PR$ Edge type in the predicted graph
$\delta$ Threshold for edge weights
$X,Y$ causal variable
$\theta,\gamma$ regression parameter
$\epsilon_{Y,i},\epsilon_{X,i}$ noise term
$\tilde{w}$ confounding-free regression parameter
$A_{i}$ exogenous variable
$H_{i}$ unobserved term
$\mathcal{P}$ class of distributions
$C$ conditioning set